# GraphAny: A Foundation Model for Node Classification on Any Graph

### v2

@@ 4.4 Ablation Study on Entropy Normalization
From the performance perspective, we observe a distinct trend in the generalization performance of normalized and unnormalized features. Figure 9 shows that unnormalized features, such as Euclidean distance and Jensen-Shannon divergence, provides a better test performance in the transductive setting. However, their test performance in the inductive setting diverges as training continues, indicating the model overfits the transductive information in the features, e.g. the scale of features. By contrast, entropy-normalized features achieve steady convergence in both transductive and inductive performance. Besides, the performance of entropy normalization is robust to different entropy values.
Figure 7: Cora (7 classes) vs Full Cora (70 classes). Left two plots: Without entropy normalization, there is a significant disparity in feature distributions between datasets. Right two plots: With entropy normalization applied, the disparity in feature distributions across datasets is largely reduced.
@@
4.5 Ablation Study on Attention Parameterization
Our attention is parameterized as a function of distance features between LinearGNN predictions on each node (node attention). Here we consider two variants of attention parameterization: (1) Graph attention uses distance features averaged over all nodes in a training batch, which results in the same attention for all nodes in a training batch. (2) Transductive attention directly learns attention weights as parameters similar to transductive methods, and assumes they can transfer to new graphs.
Figure 9 plots the test performance curves for different attention parameterizations. We notice that transductive attention does not even learn anything useful, given its performance is worse than a single LinearSGC2 model (cf. Table 2). Comparing node attention and graph attention, we can see that node attention converges slower than graph attention, but results in better performance in both transductive and inductive settings. This suggests the effectiveness of learning fine-grained attention for each node to fuse multiple LinearGNNs in GraphAny.
@@ 5 Conclusion
Limitations and Future Work. Although GraphAny showcases the capacity of graph foundation models for node classification on any graph, there are several limitations and open questions. First, LinearGNNs have limited expressiveness and cannot fit complex functions on graphs, which may restrict the overall expressiveness of GraphAny even if we combine LinearGNNs with learned attention for each node. Second, GraphAny can only solve node classification tasks at the moment, excluding regression tasks, edge-level or graph-level tasks. Besides, it remains an open question how to apply GraphAny to relational graphs, since non-pararmetric models have never been shown effective for relational graphs. We plan to address these open questions in our future work.
Societal Impact. GraphAny provides a efficient and effective tool for node classification on any graph. On the positive side, this largely reduces the carbon footprint of developing machine learning models on graphs. On a slightly negative side, our tool makes node classification more accessible and users can easily apply our tool to domains other than those studied in this paper.
Conclusion. This paper presents GraphAny, a foundation model for node classification on any graph with any feature and label spaces. GraphAny consists of two main components, LinearGNNs and an inductive attention module. LinearGNNs efficiently solve inductive inference on new graphs, while the inductive attention module learns an attention to combine multiple LinearGNN predictions. By training on a single dataset, GraphAny can effectively generalize to 30 new graphs and even surpass the average performance of transductive models supervisedly trained on each dataset.
@@ A.1 Label-permutation Invariance
In this section, we show the label-permutation invariance of the distance of LinearGNN predictions. I.e. for any two LinearGNNs $i$ and $j$, the (squared) distances between the permuted predictions, denoted as $\|\tilde{{\bm{y}}}_{u}^{(i)}-\tilde{{\bm{y}}}_{u}^{(j)})\|^{2}$ and the distances between the original predictions, i.e. $\|\hat{{\bm{y}}}_{u}^{(i)}-\hat{{\bm{y}}}_{u}^{(j)}\|^{2}$ are the same.
Consider permuting the label dimension, i.e. right multiplication with a permutation matrix ${\bm{Q}}\in\mathbb{R}^{c\times c}$. The permutated label logits $\tilde{{\bm{Y}}}$ can be expressed as:
@@
Thus, label-permutation equivariance is proved. As the distances between the permutation-equivariant distributions remain invariant, we have:
@@ A.2 Feature-permutation Invariance
Here, ${\bm{P}}{\bm{P}}^{\top}={\bm{I}}$ where ${\bm{I}}\in{\mathbb{R}}^{d\times d}$ is the identity matrix, hence proving that the predictions with feature-permutation are identical to the original predictions $\hat{{\bm{Y}}}$. That is, the predictions are feature-permutation invariant.
@@ Appendix B Datasets
We collect a diverse collection of 31 node classification datasets from three sources: PyG [12], DGL [43], and OGB [19]. We follow the default split if there is a given one, otherwise, we use the standard semi-supervised setting [21] where 20 nodes are randomly selected as training nodes for each label. The detailed dataset information is summarized in Table 3.
@@
Dataset Nodes Edges Feature Dim Num classes Num labeled nodes Category Source
Air Brazil 131 1074 131 4 80 Homophilic [32]
Cornell 183 298 1703 5 87 Heterophilic [28]
Texas 183 558 1703 5 87 Heterophilic [28]
Wisconsin 251 515 1703 5 120 Heterophilic [28]
Air EU 399 5995 399 4 80 Homophilic [32]
Air US 1190 13599 1190 4 80 Homophilic [32]
Chameleon 2277 36101 2325 5 1092 Heterophilic [28]
Wiki 2405 17981 4973 17 340 Homophilic [46]
Cora 2708 10556 1433 7 140 Homophilic [47]
Citeseer 3327 9104 3703 6 120 Homophilic [47]
BlogCatalog 5196 343486 8189 6 120 Homophilic [46]
Squirrel 5201 217073 2089 5 2496 Heterophilic [28]
Actor 7600 30019 932 5 3648 Heterophilic [28]
LastFM Asia 7624 55612 128 18 360 Homophilic [33]
AmzPhoto 7650 238162 745 8 160 Homophilic [35]
Minesweeper 10000 78804 7 2 5000 Heterophilic [30]
WikiCS 11701 431726 300 10 580 Homophilic [27]
Tolokers 11758 1038000 10 2 5879 Heterophilic [30]
AmzComp 13752 491722 767 10 200 Homophilic [35]
DBLP 17716 105734 1639 4 80 Homophilic [6]
CoCS 18333 163788 6805 15 300 Homophilic [35]
Pubmed 19717 88648 500 3 60 Homophilic [47]
FullCora 19793 126842 8710 70 1400 Homophilic [6]
Roman Empire 22662 65854 300 18 11331 Heterophilic [30]
Amazon Ratings 24492 186100 300 5 12246 Heterophilic [30]
Deezer 28281 185504 128 2 40 Homophilic [33]
CoPhysics 34493 495924 8415 5 100 Homophilic [35]
Questions 48921 307080 301 2 24460 Heterophilic [30]
Arxiv 169343 1166243 128 40 90941 Homophilic [19]
Reddit 232965 114615892 602 41 153431 Homophilic [15]
Product 2449029 123718280 100 47 196615 Homophilic [19]
@@ C.1 GraphAny Implementation
Dataset # Batches Learning Rate Hidden Dimension # MLP Layers Entropy
Cora 500 0.0002 64 1 2
Wisconsin 1000 0.0002 32 2 1
Arxiv 1000 0.0002 128 2 1
Products 1000 0.0002 128 2 1
@@ C.2 Baseline Implementation
We utilize the Deep Graph Library (DGL) [43] implementations of GCN and GAT for our baseline models. To optimize their performance, we conducted a comprehensive hyperparameter tuning using a grid search strategy on each dataset. The search space included the following parameters: number of epochs fixed at 400; hidden dimensions explored were 64, 128, and 256; number of layers tested included 1, 2, and 3; and the learning rates considered were 0.0002, 0.0005, 0.001, and 0.002.
@@ Appendix D Complete Results
Dataset MLP GCN GAT GraphAny GraphAny GraphAny GraphAny (Products) (Arxiv) (Wisconsin) (Cora) Actor 35.33$\pm$0.64 26.05$\pm$0.23 27.82$\pm$0.28 28.99$\pm$0.61 28.60$\pm$0.21 29.51$\pm$0.55 27.91$\pm$0.16 AirBrazil 25.39$\pm$5.16 40.00$\pm$2.11 47.69$\pm$7.50 34.61$\pm$16.54 34.61$\pm$16.09 36.15$\pm$16.68 33.07$\pm$16.68 AirEU 25.00$\pm$4.92 36.00$\pm$0.56 32.50$\pm$7.41 41.75$\pm$6.84 41.50$\pm$6.50 41.13$\pm$6.02 40.50$\pm$7.01 AirUS 24.90$\pm$1.14 44.21$\pm$0.77 46.60$\pm$2.49 43.57$\pm$2.07 43.64$\pm$1.83 43.86$\pm$1.44 43.46$\pm$1.45 AmzComp 59.53$\pm$1.26 85.05$\pm$0.35 85.11$\pm$0.84 82.90$\pm$1.25 83.04$\pm$1.24 82.00$\pm$1.14 82.99$\pm$1.22 AmzPhoto 73.41$\pm$2.08 90.86$\pm$0.56 91.31$\pm$0.44 90.64$\pm$0.82 90.60$\pm$0.82 90.18$\pm$0.91 90.14$\pm$0.93 AmzRatings 48.99$\pm$0.67 47.71$\pm$0.22 46.70$\pm$0.21 42.70$\pm$0.10 42.74$\pm$0.12 42.57$\pm$0.34 42.84$\pm$0.04 ogbn-arxiv 54.87$\pm$0.15 69.87$\pm$0.18 69.75$\pm$0.17 58.58$\pm$0.11 58.68$\pm$0.17 57.79$\pm$0.56 58.62$\pm$0.05 BlogCatalog 61.47$\pm$0.25 71.46$\pm$0.32 64.93$\pm$1.81 74.73$\pm$3.19 73.63$\pm$2.95 77.69$\pm$1.90 72.52$\pm$3.22 Chameleon 37.72$\pm$1.45 65.88$\pm$1.11 65.53$\pm$1.81 62.59$\pm$0.87 62.59$\pm$0.86 60.09$\pm$1.93 61.49$\pm$1.88 Citeseer 48.56$\pm$0.27 64.52$\pm$0.89 68.72$\pm$1.01 67.94$\pm$0.29 68.34$\pm$0.23 67.50$\pm$0.44 68.90$\pm$0.07 CoCS 85.20$\pm$0.10 90.72$\pm$0.05 88.84$\pm$0.31 90.46$\pm$0.54 90.45$\pm$0.59 90.85$\pm$0.63 90.47$\pm$0.63 CoPhysics 82.57$\pm$0.08 92.96$\pm$0.05 92.21$\pm$0.39 92.66$\pm$0.52 92.69$\pm$0.52 92.54$\pm$0.43 92.70$\pm$0.54 Cora 48.42$\pm$0.63 78.36$\pm$0.28 78.64$\pm$0.21 79.36$\pm$0.23 79.38$\pm$0.16 77.82$\pm$1.15 80.18$\pm$0.13 Cornell 65.40$\pm$1.21 30.27$\pm$2.26 38.92$\pm$1.48 64.86$\pm$0.00 65.94$\pm$1.48 66.49$\pm$1.48 64.86$\pm$1.91 DBLP 55.47$\pm$0.18 67.90$\pm$0.24 71.78$\pm$0.48 70.62$\pm$0.97 70.90$\pm$0.88 70.13$\pm$0.77 71.73$\pm$0.94 Deezer 54.78$\pm$0.42 54.22$\pm$0.27 54.71$\pm$1.06 52.09$\pm$2.78 52.11$\pm$2.79 52.13$\pm$3.02 51.98$\pm$2.79 FullCora 34.12$\pm$0.53 61.46$\pm$0.07 60.31$\pm$0.14 57.13$\pm$0.37 57.25$\pm$0.43 56.29$\pm$0.17 56.73$\pm$0.41 LastFMAsia 57.51$\pm$0.16 81.85$\pm$0.38 83.15$\pm$0.05 80.17$\pm$0.44 80.60$\pm$0.58 79.47$\pm$1.23 80.83$\pm$0.41 Minesweeper 80.00$\pm$0.00 80.64$\pm$0.19 80.13$\pm$0.05 80.27$\pm$0.16 80.30$\pm$0.13 80.13$\pm$0.09 80.46$\pm$0.15 ogbn-products 58.54$\pm$0.04 75.79$\pm$0.12 36.47$\pm$3.09 61.19$\pm$0.23 61.31$\pm$0.20 60.28$\pm$0.80 61.60$\pm$0.10 Pubmed 67.06$\pm$1.81 76.32$\pm$0.37 77.06$\pm$0.51 76.54$\pm$0.34 76.36$\pm$0.17 77.46$\pm$0.30 76.60$\pm$0.31 Questions 97.23$\pm$0.02 97.07$\pm$0.01 97.06$\pm$0.02 97.10$\pm$0.01 97.09$\pm$0.02 97.11$\pm$0.00 97.06$\pm$0.03 Reddit 72.17$\pm$0.11 94.87$\pm$0.02 92.43$\pm$0.48 90.67$\pm$0.13 90.58$\pm$0.12 91.00$\pm$0.24 90.46$\pm$0.03 RomanEmpire 66.35$\pm$0.28 44.91$\pm$0.24 43.46$\pm$1.23 64.66$\pm$0.84 64.25$\pm$1.09 64.06$\pm$0.78 64.25$\pm$0.64 Squirrel 31.55$\pm$0.44 47.99$\pm$0.22 45.65$\pm$1.61 49.45$\pm$0.67 49.70$\pm$0.95 42.34$\pm$3.46 48.49$\pm$0.98 Texas 58.92$\pm$6.73 52.97$\pm$2.42 60.54$\pm$6.22 73.52$\pm$2.96 72.97$\pm$2.71 73.51$\pm$1.21 71.89$\pm$1.48 Tolokers 78.16$\pm$0.00 79.63$\pm$0.58 78.41$\pm$0.22 78.18$\pm$0.03 78.18$\pm$0.04 78.24$\pm$0.03 78.20$\pm$0.02 Wiki 68.71$\pm$0.66 73.11$\pm$0.97 53.81$\pm$4.02 63.08$\pm$3.61 62.96$\pm$3.68 61.10$\pm$4.36 60.56$\pm$3.62 Wisconsin 64.71$\pm$2.78 53.72$\pm$1.07 44.31$\pm$8.16 65.89$\pm$2.23 65.10$\pm$3.22 71.77$\pm$5.98 61.18$\pm$5.08 WikiCS 72.78$\pm$0.09 79.11$\pm$0.14 79.68$\pm$0.22 75.01$\pm$0.54 74.95$\pm$0.61 73.77$\pm$0.83 74.39$\pm$0.71

### v3

@@ 4.4 Ablation Studies
Entropy Normalization. A key component of GraphAny is the entropy normalization technique applied to distance features. As discussed in Figure 5, entropy normalization ensures that the generated features are of a consistent scale. Here, we further evaluate entropy normalization from a performance perspective. Figure 9 demonstrates that unnormalized features, such as Euclidean distance and Jensen-Shannon divergence, yield better test performance in the transductive setting. However, their performance in the inductive setting decreases as training progresses, indicating that the model overfits to transductive information in the features, such as feature scale. In contrast, distance features normalized to the same entropy $H$ (denoted as EntNorm-${H}$ in the figure) achieve stable convergence in both transductive and inductive settings. Additionally, entropy normalization shows robustness to the choice of the hyperparameter entropy value $H$, with different selections resulting in similar convergence rates and performance.
@@
Attention Parameterization. It is known that the optimal message-passing patterns vary for different graphs (Zhu et al., 2020) or even different nodes (Luan et al., 2022). Therefore, GraphAny utilizes a node-level attention to adaptively combine different LinearGNN predictions. Here we consider two variants of attention parameterization: (1) Graph-level attention uses distance features averaged over all nodes in a training batch, which results in the same attention for all nodes in a training batch, losing the personalization for each node. (2) Transductive attention directly parameterizes attention weights as a $t$-dimensional vector and assumes they can transfer to new graphs. Figure 9 plots the transductive and inductive test performance curves for different attention parameterizations. We notice that transductive attention does not even learn anything useful, given its performance is worse than a single LinearSGC2 model (see Table 2). Comparing node-level attention and graph-level attention, we can see that node-level attention converges slower than graph-level attention, but results in better performance in both transductive and inductive settings. This suggests the effectiveness of learning fine-grained attention based on the local information of each node.
@@ 5 Conclusion
In this paper, we propose GraphAny, the first fully-inductive node classification model capable of performing inference on any graph with arbitrary feature or label space. GraphAny is composed of two core components: LinearGNNs and an inductive attention module. LinearGNNs enable efficient inductive inference on unseen graphs, while the inductive attention module learns to adaptively aggregate predictions from multiple LinearGNNs. Trained on a single graph, GraphAny demonstrates strong generalization to 30 new graphs, even surpassing the average performance of transductive models that are trained separately on each dataset.
Acknoledgement
The authors would like to thank Dinghuai Zhang for his helpful discussions and comments. This project is supported by the Intel-MILA partnership program, the Natural Sciences and Engineering Research Council (NSERC) Discovery Grant, the Canada CIFAR AI Chair Program, collaboration grants between Microsoft Research and Mila, Tencent AI Lab Rhino-Bird Gift Fund and a NRC Collaborative R&D Project (AI4DCORE-06). This project was also partially funded by IVADO Fundamental Research Project grant PRF-2019-3583139727. M.B. is partially supported by the EPSRC Turing AI World-Leading Research Fellowship No. EP/X040062/1 and EPSRC AI Hub No. EP/Y028872/1. The computation resource of this project is supported by Mila, Calcul Québec and the Digital Research Alliance of Canada.
@@ Appendix A Proof of Feature and Label Permutation Invariance
In this section, we prove that the proposed distance-based features have the desiring property of feature- and label-permutation invariance.
@@ A.1 Label-permutation Invariance
Here, we show the label-permutation invariance of the distance of LinearGNN predictions. I.e. for any two LinearGNNs $i$ and $j$, the (squared) distances between the permuted predictions, denoted as $\|\tilde{{\bm{y}}}_{u}^{(i)}-\tilde{{\bm{y}}}_{u}^{(j)})\|^{2}$ and the distances between the original predictions, i.e. $\|\hat{{\bm{y}}}_{u}^{(i)}-\hat{{\bm{y}}}_{u}^{(j)}\|^{2}$ are the same.
Consider permuting the label dimension, i.e. right multiplicate a permutation matrix ${\bm{Q}}\in\mathbb{R}^{c\times c}$ to the LinearGNN prediction $\hat{{\bm{Y}}}={\bm{F}}{\bm{F}}_{L}^{+}{\bm{Y}}_{L}.$ (Eq 4). The permutated label logits $\tilde{{\bm{Y}}}$ can be expressed as:
@@
Thus, label-permutation equivariance is proved. Since the permutation matrix $\mathbf{Q}$ is orthonormal (i.e. $\mathbf{Q}^{T}\mathbf{Q}=\mathbf{I}$), pairwise distances between permuted logits remain the same as those between the original logits.
@@ A.2 Feature-permutation Invariance
Here, ${\bm{P}}{\bm{P}}^{\top}={\bm{I}}$ where ${\bm{I}}\in{\mathbb{R}}^{d\times d}$ is the identity matrix, hence proving that the predictions with feature-permutation are identical to the original predictions $\hat{{\bm{Y}}}$. That is, the predictions are feature-permutation invariant. Therefore, the distance features are also feature-permutation invariant.
@@ Appendix B Datasets
We collect a diverse collection of 31 node classification datasets from three sources: PyG (Fey and Lenssen, 2019), DGL (Wang et al., 2020), and OGB (Hu et al., 2021b). We follow the default split if there is a given one, otherwise, we use the standard semi-supervised setting (Kipf and Welling, 2017) where 20 nodes are randomly selected as training nodes for each label. The detailed dataset information is summarized in Table 3.
@@
Dataset #Nodes #Edges #Feature #Classes #Labeled Nodes Train/Val/Test Ratios (%) Category Source Air Brazil 131 1074 131 4 80 61.1/19.1/19.8 Homophilic Ribeiro et al. (2017) Cornell 183 554 1703 5 87 47.5/32.2/20.2 Heterophilic Pei et al. (2020) Texas 183 558 1703 5 87 47.5/31.7/20.2 Heterophilic Pei et al. (2020) Wisconsin 251 900 1703 5 120 47.8/31.9/20.3 Heterophilic Pei et al. (2020) Air EU 399 5995 399 4 80 20.1/39.8/40.1 Homophilic Ribeiro et al. (2017) Air US 1190 13599 1190 4 80 6.7/46.6/46.6 Homophilic Ribeiro et al. (2017) Chameleon 2277 36101 2325 5 1092 48.0/32.0/20.0 Heterophilic Pei et al. (2020) Wiki 2405 17981 4973 17 340 14.1/42.9/43.0 Homophilic Yang et al. (2023) Cora 2708 10556 1433 7 140 5.2/18.5/36.9 Homophilic Yang et al. (2016) Citeseer 3327 9104 3703 6 120 3.6/15.0/30.1 Homophilic Yang et al. (2016) BlogCatalog 5196 343486 8189 6 120 2.3/48.8/48.8 Homophilic Yang et al. (2023) Squirrel 5201 217073 2089 5 2496 48.0/32.0/20.0 Heterophilic Pei et al. (2020) Actor 7600 30019 932 5 3648 48.0/32.0/20.0 Heterophilic Pei et al. (2020) LastFM Asia 7624 55612 128 18 360 4.7/47.6/47.6 Homophilic Rozemberczki et al. (2021) AmzPhoto 7650 238162 745 8 160 2.1/49.0/49.0 Homophilic Shchur et al. (2018) Minesweeper 10000 78804 7 2 5000 50.0/25.0/25.0 Heterophilic Platonov et al. (2023) WikiCS 11701 431206 300 10 580 5.0/15.1/49.9 Homophilic Mernyei and Cangea (2020) Tolokers 11758 1038000 10 2 5879 50.0/25.0/25.0 Heterophilic Platonov et al. (2023) AmzComp 13752 491722 767 10 200 1.5/49.3/49.3 Homophilic Shchur et al. (2018) DBLP 17716 105734 1639 4 80 0.5/49.8/49.8 Homophilic Bojchevski and Günnemann (2018) CoCS 18333 163788 6805 15 300 1.6/49.2/49.2 Homophilic Shchur et al. (2018) Pubmed 19717 88648 500 3 60 0.3/2.5/5.1 Homophilic Yang et al. (2016) FullCora 19793 126842 8710 70 1400 7.1/46.5/46.5 Homophilic Bojchevski and Günnemann (2018) Roman Empire 22662 65854 300 18 11331 50.0/25.0/25.0 Heterophilic Platonov et al. (2023) Amazon Ratings 24492 186100 300 5 12246 50.0/25.0/25.0 Heterophilic Platonov et al. (2023) Deezer 28281 185504 128 2 40 0.1/49.9/49.9 Homophilic Rozemberczki et al. (2021) CoPhysics 34493 495924 8415 5 100 0.3/49.9/49.9 Homophilic Shchur et al. (2018) Questions 48921 307080 301 2 24460 50.0/25.0/25.0 Heterophilic Platonov et al. (2023) Arxiv 169343 1166243 128 40 90941 53.7/17.6/28.7 Homophilic Hu et al. (2021b) Reddit 232965 114615892 602 41 153431 65.9/10.2/23.9 Homophilic Hamilton et al. (2017) Product 2449029 123718280 100 47 196615 8.0/1.6/90.4 Homophilic Hu et al. (2021b)
@@ C.1 GraphAny Implementation
Dataset # Batches Learning Rate Hidden Dimension # MLP Layers Entropy Cora 500 0.0002 64 1 2 Wisconsin 1000 0.0002 32 2 1 Arxiv 1000 0.0002 128 2 1 Products 1000 0.0002 128 2 1
@@ C.2 Baseline Implementation
We utilize the Deep Graph Library (DGL) (Wang et al., 2020) implementations of GCN and GAT for our baseline models. To optimize their performance, we conducted a comprehensive hyperparameter tuning using a grid search strategy on each dataset. The search space included the following parameters: number of epochs fixed at 400; hidden dimensions explored were 64, 128, and 256; number of layers tested included 1, 2, and 3; and the learning rates considered were 0.0002, 0.0005, 0.001, and 0.002.
@@ Appendix D Complete Results
Dataset MLP GCN GAT GraphAny GraphAny GraphAny GraphAny (Products) (Arxiv) (Wisconsin) (Cora) Actor 33.95$\pm$0.80 28.55$\pm$0.68 27.30$\pm$0.22 28.99$\pm$0.61 28.60$\pm$0.21 29.51$\pm$0.55 27.91$\pm$0.16 AirBrazil 23.08$\pm$5.83 42.31$\pm$7.98 57.69$\pm$14.75 34.61$\pm$16.54 34.61$\pm$16.09 36.15$\pm$16.68 33.07$\pm$16.68 AirEU 21.25$\pm$2.31 41.88$\pm$3.60 32.50$\pm$8.45 41.75$\pm$6.84 41.50$\pm$6.50 41.13$\pm$6.02 40.50$\pm$7.01 AirUS 22.88$\pm$1.46 46.49$\pm$1.81 48.47$\pm$4.17 43.57$\pm$2.07 43.64$\pm$1.83 43.86$\pm$1.44 43.46$\pm$1.45 AmzComp 58.28$\pm$2.98 85.83$\pm$0.86 87.01$\pm$0.50 82.90$\pm$1.25 83.04$\pm$1.24 82.00$\pm$1.14 82.99$\pm$1.22 AmzPhoto 68.20$\pm$0.88 91.88$\pm$0.79 91.86$\pm$1.07 90.64$\pm$0.82 90.60$\pm$0.82 90.18$\pm$0.91 90.14$\pm$0.93 AmzRatings 47.90$\pm$0.45 47.35$\pm$0.26 47.18$\pm$0.42 42.70$\pm$0.10 42.74$\pm$0.12 42.57$\pm$0.34 42.84$\pm$0.04 BlogCatalog 64.11$\pm$1.95 71.51$\pm$2.62 61.98$\pm$5.99 74.73$\pm$3.19 73.63$\pm$2.95 77.69$\pm$1.90 72.52$\pm$3.22 Chameleon 36.62$\pm$0.87 64.69$\pm$2.21 67.76$\pm$0.72 62.59$\pm$0.87 62.59$\pm$0.86 60.09$\pm$1.93 61.49$\pm$1.88 Citeseer 44.40$\pm$0.44 63.40$\pm$0.63 69.10$\pm$1.59 67.94$\pm$0.29 68.34$\pm$0.23 67.50$\pm$0.44 68.90$\pm$0.07 CoCS 85.88$\pm$0.93 91.83$\pm$0.71 88.47$\pm$0.79 90.46$\pm$0.54 90.45$\pm$0.59 90.85$\pm$0.63 90.47$\pm$0.63 CoPhysics 87.43$\pm$1.98 93.93$\pm$0.37 93.01$\pm$0.89 92.66$\pm$0.52 92.69$\pm$0.52 92.54$\pm$0.43 92.70$\pm$0.54 Cora 48.42$\pm$0.63 81.40$\pm$0.70 81.70$\pm$1.43 79.36$\pm$0.23 79.38$\pm$0.16 77.82$\pm$1.15 80.18$\pm$0.13 Cornell 67.57$\pm$5.06 35.14$\pm$6.51 35.14$\pm$3.52 64.86$\pm$0.00 65.94$\pm$1.48 66.49$\pm$1.48 64.86$\pm$1.91 DBLP 56.27$\pm$0.62 73.02$\pm$2.22 73.87$\pm$1.35 70.62$\pm$0.97 70.90$\pm$0.88 70.13$\pm$0.77 71.73$\pm$0.94 Deezer 54.24$\pm$2.15 53.69$\pm$2.29 55.99$\pm$3.78 52.09$\pm$2.78 52.11$\pm$2.79 52.13$\pm$3.02 51.98$\pm$2.79 LastFMAsia 57.41$\pm$0.93 81.91$\pm$1.12 84.66$\pm$0.61 80.17$\pm$0.44 80.60$\pm$0.58 79.47$\pm$1.23 80.83$\pm$0.41 Minesweeper 80.00$\pm$0.00 81.12$\pm$0.37 80.08$\pm$0.04 80.27$\pm$0.16 80.30$\pm$0.13 80.13$\pm$0.09 80.46$\pm$0.15 Pubmed 69.50$\pm$1.79 76.60$\pm$0.32 77.30$\pm$0.60 76.54$\pm$0.34 76.36$\pm$0.17 77.46$\pm$0.30 76.60$\pm$0.31 Questions 97.33$\pm$0.06 97.15$\pm$0.04 97.11$\pm$0.02 97.10$\pm$0.01 97.09$\pm$0.02 97.11$\pm$0.00 97.06$\pm$0.03 Reddit 72.16$\pm$0.15 94.89$\pm$0.02 92.76$\pm$0.46 90.67$\pm$0.13 90.58$\pm$0.12 91.00$\pm$0.24 90.46$\pm$0.03 Roman 65.80$\pm$0.35 45.08$\pm$0.43 43.93$\pm$0.45 64.66$\pm$0.84 64.25$\pm$1.09 64.06$\pm$0.78 64.25$\pm$0.64 Squirrel 30.36$\pm$0.78 47.07$\pm$0.71 46.69$\pm$1.44 49.45$\pm$0.67 49.70$\pm$0.95 42.34$\pm$3.46 48.49$\pm$0.98 Texas 48.65$\pm$4.01 51.35$\pm$2.71 54.05$\pm$2.41 73.52$\pm$2.96 72.97$\pm$2.71 73.51$\pm$1.21 71.89$\pm$1.48 Tolokers 78.16$\pm$0.02 79.93$\pm$0.10 78.50$\pm$0.55 78.18$\pm$0.03 78.18$\pm$0.04 78.24$\pm$0.03 78.20$\pm$0.02 Wiki 63.79$\pm$1.77 70.09$\pm$1.51 59.63$\pm$4.16 63.08$\pm$3.61 62.96$\pm$3.68 61.10$\pm$4.36 60.56$\pm$3.62 Wisconsin 66.67$\pm$3.51 37.25$\pm$1.64 52.94$\pm$3.10 65.89$\pm$2.23 65.10$\pm$3.22 71.77$\pm$5.98 61.18$\pm$5.08 WikiCS 72.72$\pm$0.43 79.12$\pm$0.45 79.27$\pm$0.20 75.01$\pm$0.54 74.95$\pm$0.61 73.77$\pm$0.83 74.39$\pm$0.71 OGBN-Arxiv 55.50$\pm$0.23 71.74$\pm$0.29 73.65$\pm$0.11 58.58$\pm$0.11 58.68$\pm$0.17 57.79$\pm$0.56 58.62$\pm$0.05 OGBN-Products 61.06$\pm$0.08 75.79$\pm$0.12 79.45$\pm$0.59 61.19$\pm$0.23 61.31$\pm$0.20 60.28$\pm$0.80 61.60$\pm$0.10 FullCora 33.54$\pm$0.64 61.06$\pm$0.24 58.95$\pm$0.55 57.13$\pm$0.37 57.25$\pm$0.43 56.29$\pm$0.17 56.73$\pm$0.41
Appendix E Additional Experimental Results
E.1 Ablation Study on More Graph Operators
In this section, we further validate the effectiveness of GraphAny by gradually adding diverse types of graph convolution for propagating features. Specifically, we consider three types of graph convolutions:
  • GraphAny-LinearGNN gradually adds the following graph convolution operators: ${\bm{F}}={\bm{X}}$ (Linear), ${\bm{F}}=({\bm{I}}-\bar{{\bm{A}}}){\bm{X}}$ (LinearHGC1), ${\bm{F}}=({\bm{I}}-\bar{{\bm{A}}})^{2}{\bm{X}}$ (LinearHGC2), ${\bm{F}}=\bar{{\bm{A}}}{\bm{X}}$ (LinearSGC1), and ${\bm{F}}=\bar{{\bm{A}}}^{2}{\bm{X}}$ (LinearSGC2), with $\bar{{\bm{A}}}$ denoting the row normalized adjaceny matrix. In this case, when increasing $t$ high-pass and low-pass predictions are gradually added.
  • GraphAny-Chebyshev gradually adds the Chebyshev filters (Defferrard et al., 2016), where $\bm{F}^{(t)}$ is computed recursively by:
    $\bm{F}^{(1)}=\bm{X},$ (16)
    $\bm{F}^{(2)}=\hat{\bm{L}}\bm{X},$
    $\bm{F}^{(t)}=2\hat{\bm{L}}\bm{F}^{(t-1)}-\bm{F}^{(t-2)},$
    and $\hat{\bm{L}}$ denotes the scaled and normalized Laplacian $\frac{2\bm{L}}{\lambda_{\max}}-\bm{I}$ with $\bm{L}=\bm{I}-\bm{D}^{-1/2}\bm{A}\bm{D}^{-1/2}$. In this case, when increasing $t$ high order predictions are gradually added.
  • GraphAny-PPR gradually adds the personalized PageRank with increasing restart ratios: {0.01, 0.25, 0.5, 0.75, 0.99}. In this case, when increasing $t$ local predictions are gradually added.
Figure 10: Ablation study with different graph operators (test accuracy in %): GraphAny-LinearGNN (left), GraphAny-Chebyshev (middle), and GraphAny-PPR (right). GraphAny is trained on the Wisconsin dataset and evaluated on all 31 datasets.
We compare the performance of GraphAny against the baseline, which averages all channels, denoted as MeanAgg. The results are shown in Figure 10, where we have the following observations:
GraphAny consistently outperforms the baseline MeanAgg by a significant margin, demonstrating effective and consistent inductive knowledge transfer across multiple datasets with different types of graph convolutions. The performance varies across graph convolution types. For LinearGNNs, low-pass graph convolutions (LinearSGC1 and LinearSGC2) significantly enhance performance. In contrast, for Chebyshev graph convolutions, adding high-order convolutions reduces performance as higher-order information introduces more noise. For personalized PageRank, adding more local channels consistently improves results.
E.2 Comparison Against ACM-GNN
To comprehensively compare GraphAny’s performance, we included ACM-GNN (Luan et al., 2022) as a baseline due to its good performance on both homophilic and heterophilic graphs. However, ACM-GNN faces scalability issues that prevent its application to large graphs. Using the authors’ implementation [^2] https://github.com/SitaoLuan/ACM-GNN/tree/main/ACM-Pytorch/, we encountered out-of-memory for a GPU of 40GB on four large datasets: Questions, Reddit, Arxiv, and Product. Therefore, we evaluated ACM-GNN on the remaining 27 graphs, with results reported in Table 6.
Our observations are as follows: Tuning ACM-GNN is highly time-consuming, requiring 672 GPU hours on 27 graphs, while GraphAny requires only 4 GPU hours, showcasing its efficiency and the advantage of inductive inference. In terms of performance, GraphAny outperforms ACM-SGC and is only slightly (1-2%) below ACM-GCN. However, this slight difference is not a significant disadvantage for GraphAny, given the unfair advantage transductive models have by leveraging the inductive validation set for parameter and hyperparameter tuning, as well as the substantial difference in runtime.
Table 6: Experimental results of ACM-GNNs. Hyperparameter search durations: ACM-GNNs (672 GPU hours), MLP/GCN/GAT (240 GPU hours total), GraphAny (denoted as Ours, 4 GPU hours).
Metric MLP GCN GAT ACM-SGC ACM-GCN Ours-Cora Ours-Wis Ours-Arxiv Ours-Product
Heldout (25 graphs) 54.87 64.19 64.01 60.91 66.92 65.11 65.23 65.56 65.56
Overall (27 graphs) 55.07 63.83 64.26 61.82 67.79 65.58 66.02 66.12 66.14